Mechanics

In the relation,
P = (a/b)e-(as/kT)
P is Pressure, s is distance, k is boltzmann constant, T is temperature. The dimensions of b is
  1. [M0L2T0]
  2. [ML2T]
  3. [ML0T-1]
  4. [M0L2T-1]

Saurabh Mishra
8 Years agoGrade

1 Answer

Khimraj
8 Years ago
power of e must be dimensionless
so unit of a = kT/s
unitof a/b must be unit of pressure then
b = kT/sP = {ML2T-2K-1][K]/[L][ML-1T-2]        =   [MoL2To]
Hope it clears. If you like answer then please approve it.